OpenStudy (anonymous):

LIMIT SIN^2(X)/X(1+COS(X) AS X GOES TO 0

OpenStudy (nikvist):

\[\lim_{x\rightarrow 0}\frac{\sin^2x}{x(1+\cos x)}\] is it correct?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

zero

OpenStudy (anonymous):

is the answer fairly sure

OpenStudy (anonymous):

its not a indeterminate form , we can just sub in x=0

OpenStudy (nikvist):

Ok, answer is 0, but this is indeterminate form 0/0

OpenStudy (anonymous):

ahh yeh , im an idiot , didnt see the x factor on the bottom differnetiate top and bottom and try again

OpenStudy (anonymous):

YES I GOT 0/0 BUT MY TEACHER MARK MY PAPER STILL WRONG WITH MY STEPS I JUST SUBSTITUTE X FOR 0

OpenStudy (dumbcow):

\[=\lim_{x \rightarrow 0}\frac{\sin x ^{2}}{x} * \lim_{x \rightarrow 0}\frac{1}{1+\cos x} = \lim_{x \rightarrow 0}2\sin x \cos x * \frac{1}{2} = 0\]

OpenStudy (anonymous):

so [ 2sinxcosx ] / [ x(-sin(x) ) + (1+cos(x)) ] 2sin(x)cos(x) / [ 1 +cos(x) -xsin(x) ]

OpenStudy (anonymous):

now when you sub x=0 , it isnt indeterminate , and we do get 0 as the final answer

OpenStudy (nikvist):

dumbcow, \[\lim_{x\rightarrow a}f(x)g(x)\neq\lim_{x\rightarrow a}f(x)\cdot\lim_{x\rightarrow a}g(x)\]

OpenStudy (anonymous):

^doesnt it? I thought I remember reading somewhere that it does

OpenStudy (dumbcow):

really? are you sure lim x^2 as x->2 is 4 limx * limx as x->2 is 2*2=4

OpenStudy (dumbcow):

no they can be separated, see below http://tutorial.math.lamar.edu/Classes/CalcI/LimitsProperties.aspx

OpenStudy (anonymous):

YES GUYS YOU CAN SEPARATE IM SURE OF THAT

OpenStudy (anonymous):

before you answer, shut the caps button off LOL!

OpenStudy (anonymous):

oh well ~